Exercise 13-03 a-b During its first year or operations, Flint Corporation had the following transactions pertaining to its common stock Jan. 10 Issued 65,500 shares for cash at $5 per share July 1 Issued 43,000 shares for cash at $10 per share ournalize the transactions, assuming that the common stock has a par value of $5 per share. (Record journal entries in the order presented in the problem. Credit account titles are automatically indented when amount is entered. Do not indent manually.) Date Account Titles and Explanation Debit Credit Jan. 10 27500 327500 July 1 SHOW LIST OF ACCOUNTS SHOW SOLUTION SHOW ANSWERournalize the transactions, assuming that the common stock is no-par with a stated value of $2 per share. (Record journal entries in the order presented in the problem. Credit account titles are automatically Indented when amount is entered. Do not indent manually.) Date Account Titles and Explanation Debit Credit LINK TO TEXT
Answer #1

(1) Par value is $5 per stock:

Date Account titles Debit Credit
Jan 10 Cash 327,500.00
Common Stock 327,500.00
July 1 Cash 430,000.00
Common Stock 215,000.00
Paid in capital in excess of par 215,000.00

Calculations: cash = 43,000 shares*$10 = $430,000. Common stock = 43,000*$5 = 215,000 and paid in capital = 43,000*(10-5) = 215,000

(2) Issue of no par stock with a stated value of $2:

Date Account titles Debit Credit
Jan 10 Cash 327,500.00
Common Stock 131,000.00
Additional paid in capital 196,500.00
July 1 Cash 430,000.00
Common stock 86,000.00
Additional paid in capital 344,000.00

Calculations: When stocks are issued at $5 then common stock =65500 shares*$2 = 86000 and additional paid in capital = 65500 shares*(5-2) = 196500

When stocks are issued at $10 then common stock amount = 43000 shares*$2 = 86000 and additional paid in capital = 43000 shares*(10-2) = 344000